UT Jul 14 - Extra Engineering Night - rigid installed on SX




Night Info

  • Sunset 19:29
  • Nautical Twilight Ends 20:31
  • Astro Twilight Ends 21:08
  • Astro Twilight Starts 03:43
  • Nautical Twilight Starts 04:20
  • Sunrise 05:22
  • Moonrise 19:57
  • Moonset 06:15
  • Moon Illumination 99% - rise and set before and after nautical twilights

Staff

  • OSA: D.G.Huerta
  • ISp: J.Schuh
  • ISA: B.Rothberg
  • Mtn Manager:B.Bingham

Issues

  • MODS2-Blue pressure issue and readout issue - may try to take data


Closed Dome

  • Closed dome MODS cals
    • Grating cals
  • Direct mode laser on/laser off (lbto@robs:~/kuhn/).
    • You would need to execute the commands "irlaser disable", "irlaser off' in the command window.
    • The scripts do not do this. Run the disabled after the enabled (that way the collimator is aligned and at the same position as for the laser on images).
      • laser_enabled_blue.obs

      • laser_disabled_off_blue.obs

      • laser_enabled_red.obs

      • laser_disabled_off_red.obs


Night Plan

  • Goal:
    • Obtain absolute flux comparison among SX w/rigid and DX w/ASM and SX w/ASM - requires photometric conditions
    • If non-photometric conditions - then relative flux comparisons using spectrophotometric stars and photometric standards
    • Use f/15 instruments (MODS, PEPSI, LUCI)
    • See IT 8600
    • Also see IT 7730
  • Starting Configuration: bino-MODS - highest priority given forecast
    • then PEPSI, then LUCI (in order of priority)
  • Short Verificaiton rigid is installed on SX
    • Successful preset with guiding/WFS'ing on AGWs
    • stars are round
  • MODS Spectrophotometric standards
    • About 1 hr each for grating, prism.
    • Scripts: ~/Service_Observing/2022A_D_E/MODS/2022Jul13/
    • Observe 2 standards - different airmasses use dual, blue, and red gratings - highest priority
      • BD+33 2642 15 51 59 +32 56 54 - twilight-01:45hr - 60,75 deg from Moon
        • BD+33 observed a few days ago in clear conditions - 20220707
      • Feige110 23 19 58 -05 09 56 - visible 01hr-twlight - 78, 60deg from Moon
    • Photometric standards (all sloan filters)
      • M13 16 41 41 +36 27 40 - visible twlight-01:45 - 60-70 deg from Moon
      • M92 17 17 07 +43 08 09 - visible twilight-03:45hr - 73-79 deg from Moon
      • M71 19 53 46 +18 46 45 - visible all night - 53-48 deg from Moon
    • Prism mode for spectrophotometric standards (if clear)
      • Feige110 23 19 58 -05 09 56 - visible 01hr-twlight
        • flux tables for both 10-A and 50-A bins.
  • PEPSI Flux Standards Available for the Night
    • HR 5501 14 45 30 +00 43 02 V=5.6 - visibile twilight-23:15hr
    • HR 7596 19 54 44 +00 16 25 V=5.6 -visible 21:15-twilight
    • HR 7950 20 47 40 -09 29 44 V=3.7 - visible 22:45-twilight
    • HR 153 00 36 58 +53 53 48 V=3.6 - visible midnight-twilight
    • Restriced blue CD setting to avoid hardware issues
  • LUCI Photometric standards
    • use J,H, and K both sides - relative or absolute flux
      • Y1 comparisons - cutoff is 0.95 micron on blue-end - matches LUCI-2 dichroic cutoff
    • OT file: /home/lbto/Service_Observing/2022A_D_E/PoorWeather/LUCI_Phot_Standards/LUCI_Phot_ZP_OT_V5.xml
    • Standards available all night
  • Other (tasks lower priority to try once above are completed)
    • MODS trailed images issue

Night Log

Closed Dome Activities

  • UT 01:28 - starting MODS GUIs, waking MODS and taking simSnaps - all OK except M2B sequencing is messed up
  • UT 01:39 - bringing up LUCI GUIs, init, test darks - OK
  • UT 01:48 - PEPSI GUI up on VNC - just wating for clouds to flitter away
  • UT 02:05 - DG.Huerta reconfiguring for bino-MODS
  • UT 02:54 - current allsky
  • AllSkyCurrentImage-1.JPG
  • UT 03:12 - test preset with MODS
  • UT 04:01 - still closed for thick clouds overhead - but so far no rain
  • UT 04:17 - lights out in dome - going to take some grating cals
    • Pixel Flats Red Only
    • Arcs Red Only
  • UT 05:29 - doing test for IMCS laser and 843 nm line
    • 2.4" LS
    • Observing mode
    • instconfig dual grating
    • 1800sec, manually engage IMCS
    • Elevation 90
    • UT 06:07 Moving to Elevation 60
      • relocking IMCS
      • M1B 0004
      • M1R 0012
      • M2B 0004 - doesn't look corrupted
      • M2R 0012
    • UT 06:43 - moving to Elevation 30
      • relocking IMCS
      • M1B 0005
      • M1R 0012
      • M2B 0005 - doesn't look corrupted
      • M2R 0012
    • UT 07:18 - moving back to zenith
    • UT 07:19 - dual imaging flats dual grating (2 exposure per band)
      • Error M1: [2022-07-14T07:20:05.772] M1.IE>MC1 ERROR: RCAMFOC RCAMFOC=BUSY
      • should have gone to 1160
        • M1B 0006-0009
        • M1R 0013-0019
        • M2B 0006-0009 -these do look corrupted
        • M2R 0013-0019
    • UT 07:38 - moving to horizon - raining
    • UT 11:10 - putting both MODS to bed

-- BarryRothberg - 13 Jul 2022
Topic revision: r6 - 14 Jul 2022, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback